# Cuipo Jiang and Jingtian Song

### v2

@@ 4.2 Associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1},m\geqslant 1$
$\begin{array}[]{ll}&v^{1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}^{2m+1}(h_{1}+h_{2})(-1)^{2m}\\ \\ &+e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
@@
$\begin{array}[]{ll}&cv^{2}-e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m+1}h_{2}^{2m}(h_{1}+h_{2})(-1)^{2m}\\ \\ &-e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
@@
$\mathfrak{g}=G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})\cup G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})\cup G\cdot({\mathfrak{h}}^{reg})\cup{\mathcal{N}},$
@@
$G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})\cup{\mathcal{N}}\subseteq\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
@@
Then by Lemmas 4.7-4.8, the zero locus of $\overline{{\mathcal{I}}_{k}}\subseteq R_{V^{k}(sl_{3})}$ is exactly $\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$ It follows that
$X_{L_{k}(sl_{3})}=\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
@@
    $\mathfrak{g}^{*},\ \{0\},\ \mathcal{N},\ \overline{{\mathcal{O}}_{min}},\ \overline{G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})},\ \overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
@@
    Let $h$ be a regular semi-simple vector of $sl_{3}$, then there is no $k\in{\mathbb{C}}$ such that $X_{L_{k}(sl_{3})}=\overline{G\cdot{\mathbb{C}}^{*}h}$.
@@
${\mathcal{V}}(I_{k})=\{t\bar{\Lambda}_{1}-\frac{2i}{3}\bar{\Lambda}_{2},t\bar{\Lambda}_{2}-\frac{2i}{3}\bar{\Lambda}_{1},t\bar{\Lambda}_{1}-(t+\frac{2i}{3}+1)\bar{\Lambda}_{2},~{}t\in{\mathbb{C}},i=0,1,\cdots,2m\}.$
@@
If $m=0$, that is $k=-1$, the conjecture is true by Proposition 5.5 of [AP08]. When $m=1$, we could verify the conjecture by computer programming.
@@ 5 Varieties of simple affine $W$-algebras $W_{k}(sl_{3},f)$
Let $\mathfrak{g}$ be a simple complex Lie algebra, $f$ a regular nilpotent element of $\mathfrak{g}$, and $k=-h^{\vee}+\frac{p}{q}$ a non-degenerate admissible number. Then the simple $W$-algebra $W_{k}(g,f)$ is rational and lisse.
@@
Let $k=-1$ and $f$ a minimal nilpotent element of $sl_{3}$. It was proved in [AKF+18] that $H^{0}_{DS,f}(L_{-1}(sl_{3}))=W_{-1}(sl_{3},f)$, and is isomorphic to the rank one Heisenberg algebra $M(1)$. So its associated variety is one-dimensional.
@@
By Theorem 8.2 of [AFP23a], $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$, are irreducible ordinary modules of $L_{k}(sl_{3})$. This means that $L_{k}(sl_{3})$ has infinitely many irreducible ordinary modules. Then by Theorem 5.4 of [AKR23], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the minimal quantum hamiltonian reductions ${\mathcal{H}}_{j_{t+1,1}\Delta_{t+1,1}}$ and ${\mathcal{H}}_{j_{1,t+1}\Delta_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3},f)$-modules, where
@@
$A_{\mu}=\mu(h_{1}-h_{2})+f_{\theta}=\left[\begin{array}[]{ccc}-\mu&0&0\\ 0&2\mu&0\\ 1&0&-\mu\end{array}\right].$
@@
Let $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$ be the irreducible ordinary modules of $L_{k}(sl_{3})$ as in the proof of Theorem 5.3, then by Theorem 5.4 of [AKR23], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the principal quantum hamiltonian reductions ${\mathcal{W}}_{h_{t+1,1}w_{t+1,1}}$ and ${\mathcal{W}}_{h_{1,t+1}w_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3})$-modules, where

### v3

@@ 4.2 Associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1},m\geqslant 1$
$\begin{array}[]{ll}&v^{1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}^{2m+1}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\\ \\ &+e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
@@
$\begin{array}[]{ll}&cv^{2}-e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m+1}h_{2}^{2m}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\\ \\ &-e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
@@
$\mathfrak{g}=G.{\mathbb{C}}^{*}(h_{1}-h_{2})\cup G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})\cup G\cdot({\mathfrak{h}}^{reg})\cup{\mathcal{N}},$
@@
$G.{\mathbb{C}}^{*}(h_{1}-h_{2})\cup{\mathcal{N}}\subseteq\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
@@
Then by Lemmas 4.7-4.8, the zero locus of $\overline{{\mathcal{I}}_{k}}\subseteq R_{V^{k}(sl_{3})}$ is exactly $\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$ It follows that
$X_{L_{k}(sl_{3})}=\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
@@
    $\mathfrak{g}^{*},\ \{0\},\ \mathcal{N},\ \overline{{\mathcal{O}}_{min}},\ \overline{G.{\mathbb{C}}^{*}(h_{1}-h_{2})},\ \overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
@@
    Let $h$ be a regular semi-simple vector of $sl_{3}$, then there is no $k\in{\mathbb{C}}$ such that $X_{L_{k}(sl_{3})}=\overline{G.{\mathbb{C}}^{*}h}$.
@@
${\mathcal{V}}(I_{k})=\{t\bar{\Lambda}_{1}-\frac{2i}{2m+1}\bar{\Lambda}_{2},t\bar{\Lambda}_{2}-\frac{2i}{2m+1}\bar{\Lambda}_{1},t\bar{\Lambda}_{1}-(t+\frac{2i}{2m+1}+1)\bar{\Lambda}_{2},~{}t\in{\mathbb{C}},i=0,1,\cdots,2m\}.$
@@
[…]
@@ 5 Varieties of simple affine $W$-algebras $W_{k}(sl_{3},f)$
Let $\mathfrak{g}$ be a simple complex Lie algebra, $f$ a regular nilpotent element of $\mathfrak{g}$, and $k=-h^{\vee}+\frac{p}{q}$ a non-degenerate admissible number. Then the simple $W$-algebra $W_{k}(\mathfrak{g},f)$ is rational and lisse.
@@
Let $k=-1$ and $f$ a minimal nilpotent element of $sl_{3}$. It was proved in [AKMP+18] that $H^{0}_{DS,f}(L_{-1}(sl_{3}))=W_{-1}(sl_{3},f)$, and is isomorphic to the rank one Heisenberg algebra $M(1)$. So its associated variety is one-dimensional.
@@
By Theorem 8.2 of [AMP23], $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$, are irreducible ordinary modules of $L_{k}(sl_{3})$. This means that $L_{k}(sl_{3})$ has infinitely many irreducible ordinary modules. Then by Theorem 5.4 of [AKR24], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the minimal quantum hamiltonian reductions ${\mathcal{H}}_{j_{t+1,1}\Delta_{t+1,1}}$ and ${\mathcal{H}}_{j_{1,t+1}\Delta_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3},f)$-modules, where
@@
$A_{\mu}=\mu(h_{1}-h_{2})+f_{\theta}=\left[\begin{array}[]{ccc}-\mu&0&0\\ 0&2\mu&0\\ 1&0&-\mu\end{array}\right]$
@@
Let $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$ be the irreducible ordinary modules of $L_{k}(sl_{3})$ as in the proof of Theorem 5.3, then by Theorem 5.4 of [AKR24], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the principal quantum hamiltonian reductions ${\mathcal{W}}_{h_{t+1,1}w_{t+1,1}}$ and ${\mathcal{W}}_{h_{1,t+1}w_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3})$-modules, where
@@
6 Declarations
  1. (1)
    This work is supported by Natural Science Foundation of China (Grant number: 12171312);
  2. (2)
    The authors have no relevant financial or non-financial interests to disclose;
  3. (3)
    All datasets supporting the analysis and conclusions of the paper are publicly available at the time of publication.